OpenStudy (shicarak):

simplify the equation please -2+4(2y-4)

OpenStudy (steve816):

-2 + 4(2y-4) You can distribute! It would turn into -2 + 8y -16

OpenStudy (steve816):

Now, do you think you can simplify from here?

OpenStudy (shicarak):

I think so thank you

OpenStudy (steve816):

\(\Huge{\color{red}{\heartsuit}\color{orange}{\bigstar}\color{yellow}{\heartsuit}\color{green}{\bigstar}\color{blue}{\heartsuit}\color{purple}{\bigstar}\color{hotpink}{\heartsuit}}\)\(\Huge\bf\color{#FF0000}W\color{#FF4900}e\color{#FF9200}l\color{#FFDB00}c\color{#FFff00}o\color{#B6ff00}m\color{#6Dff00}e\color{#24ff00}~\color{#00ff00}t\color{#00ff49}o\color{#00ff92}~\color{#00ffDB}O\color{#00ffff}p\color{#00DBff}e\color{#0092ff}n\color{#0049ff}S\color{#0000ff}t\color{#2400ff}u\color{#6D00ff}d\color{#B600ff}y\color{#FF00ff}!\)\(~~~~~~~~~~~~~~~~~~~~~~~~~~\Huge{\color{red}{\heartsuit}\color{orange}{\bigstar}\color{yellow}{\heartsuit}\color{green}{\bigstar}\color{blue}{\heartsuit}\color{purple}{\bigstar}\color{hotpink}{\heartsuit}}\)\(\Large\bf\color{Lime}{Please,\ don't\ hesitate\ to\ ask\ questions!}\)\(\Large\bf\color{magenta}{We're\ here\ to\ help!}\Large\color{magenta}{\ddot\smile}\)

KjSaif (kjsaif):

Distribute: =−2+(4)(2y)+(4)(−4) =−2+8y+−16 Combine Like Terms: =−2+8y+−16 =(8y)+(−2+−16) =8y+−18

KjSaif (kjsaif):

Answer: =8y−18

OpenStudy (shicarak):

thank you KJSaif

KjSaif (kjsaif):

no problem!

OpenStudy (shicarak):

I have another one please help \[15-2(3x+5y)\]

KjSaif (kjsaif):

ask this in another question or this would be considered spam :)

OpenStudy (steve816):

Wat a spam??? No way

OpenStudy (shicarak):

how

KjSaif (kjsaif):

A mod told me that

OpenStudy (steve816):

Well, must've been an ignorant one.